Lazorwolf64:

Help me pls Mrs. Flores is planning on giving her students an ice cream cone as a reward. She plans on filling the cones and adding a 1 unit scoop on top of the cone. If she buys 553 cubic units of ice cream to share between herself and her 99 students and knows that the cones will have a height of 3 units; what is the maximum radius that the cones can be? Round to the nearest tenth and remember words deserve words. Use the pi key for calculations.

surjithayer:

99+1=100 volume of icecreams filled in 100 cones=553-100=453 volume of each cone\[=\frac{ 453 }{ 100 }=4.53\] h= 3cm \[\frac{ 1 }{ 3 } \pi r^2 \times 3=4.53\] \[r^2=\frac{ 4.53 }{ \pi } \] \[r=\sqrt{\frac{ 4.53 }{ \pi}} \approx ~1.20~cm\]

Lazorwolf64:

You are a life saver!!!!!!!

surjithayer:

yw